OpenStudy (anonymous):

What is the solution to the system. 3x + 5y = 7.60 x + 2y = 2.90

OpenStudy (anonymous):

@johnweldon1993 And here is my question!

OpenStudy (gibbs):

x + 2y = 2.90<br />=&gt; x = 2.90 - 2y<br /><br />Substitue this value of x in the other equation:<br />3x + 5y = 7.60<br />3*(2.90 - 2y) + 5y = 7.60<br />8.70 - 6y + 5y = 7.60<br />1.10 = y<br /><br />and then substitute this value of y in either equation to find x:<br />x = 2.90 - 2y = 2.90 - 2*1.10 = 2.90 - 2.20 = 0.70<br /><br />Solution: (x, y) = (0.70, 1.10)

OpenStudy (godlovesme):

help ? @BlossomCake

OpenStudy (anonymous):

Solve the following: \[\{30 x+50 y=76,10 x+20 y=29\} \]

OpenStudy (anonymous):

\[\left\{x=\frac{7}{10},y=\frac{11}{10}\right\} \]